# TC-GS: A Faster Gaussian Splatting Module Utilizing Tensor Cores

### v1

@@ 5.4.2 Training
6 Conclusion
In this paper, we propose TC-GS, a high-performance rendering module for accelerating 3D Gaussian Splatting. TC-GS optimizes the rendering pipeline by decomposing alpha-blending into two components, i.e., Gaussian-related and pixel-related matrices. Through reformulating alpha-blending as matrix multiplications, TC-GS achieves a significant speedup of 2.03$\times$ to 4.76$\times$ for alpha blending on Tensor Cores. When integrated into state-of-the-art pipelines, TC-GS attains approximately a 2$\times$ throughput improvement; even when compared against highly optimized, software-pipelined methods, it delivers a 1.38$\times$ speed-up. These gains are realized without any degradation in image quality, demonstrating TC-GS’s superior performance, scalability, and broad applicability for future real-time neural rendering systems.
Appendix A Notations
@@
$c$ Color of 3D Gaussian, only in Section 3 $W$ Matrix projecting 3D position into 2D plane
$J$ Jacobian of projection $\mu^{\prime}$ Mean of projected Gaussian
$\Sigma^{\prime}$ Covariance of projected Gaussian $G^{\prime}$ Projected Gaussian
$\mathcal{G}^{\prime}$ Set of projected Gaussians $p$ Pixel
@@
$C$ Rendered pixel color $\alpha$ Opacity of fragments.
@@
$N$ Number of splats. $\mathcal{E}$ Complexity of the pipeline in 3DGS
@@
$\begin{bmatrix}a&b\\ b&c\end{bmatrix}$ $\Sigma^{\prime}$, in Section 4 and Appendix B $(\mu^{\prime}_{x},\mu^{\prime}_{y})$ Coordinates of project mean $\mu^{\prime}$
@@
$\Delta p$ Pixel in local coordinates $\Delta\mu^{\prime}$ Projected mean in local coordinates
$r$ Arbitrary real number $H(r)$ $r$ in half-precision presentation.
$\epsilon_{H}$ Machine epsilon of half precision $\epsilon_{F}$ Machine epsilon of single precision
$\hat{\beta}$ Computed $\beta$ with TC-GS. $S$ The upper bound of the rounding error of $\beta$
$S_{1}$ The rounding error from $\mathbf{u}$ $S_{2}$ The rounding error from $\mathbf{v}$
$\lambda_{1},\lambda_{2}$ Eigenvalues of $\Sigma^{\prime}$ $\mathbf{v}_{\Delta}$ Gaussian vector in local coordinates
Table 6: Symbols used in this paper.
Appendix B Derivation of the Upper Bound of Rounding Error on TC-GS
We only consider the exponent $\beta$ in fragment which is not culled, otherwise it will not contribute the pixel color. Besides, we suppose that if a Gaussian will be rendered, each axis of its ellipse is greater than $0.5$.
B.1 Rounding Error on Global Coordinates
For an arbitrary real number $r\in\mathbb{R}$ with half-precision point representation $H(r)$, the rounding error can be the estimated by
$|H(r)-r|\leq|r|\epsilon_{H},$ (20)
where $\epsilon_{H}\approx 9.77\times 10^{-4}$ is the machine epsilon of FP16 or TF32. We can
In Tensor Cores, the products of two half-precision floating-point numbers $H(r_{1}),H(r_{2})$ is stored as FP32. Therefore, the rounding error of the product satisfies that
$|H(r_{1})H(r_{2})-r_{1}r_{2}|\leq|r_{1}r_{2}|\epsilon_{F}+|r_{1}|\epsilon_{H}+|r_{2}|\epsilon_{H}+O(\epsilon_{H}^{2}),$ (21)
where $\epsilon_{F}$ is the machine error of FP32. To simply our analysis, we ignore the smaller machine error $\epsilon_{F}$ and quadratic term $\epsilon_{H}^{2}$. Therefore, the rounding error of the product satisfies that
$|H(r_{1})H(r_{2})-r_{1}r_{2}|\leq(|r_{1}|+|r_{2}|)\epsilon_{H}$ (22)
Then, we delve into the vector product performed by Tensor Cores. Consider an arbitrary vectors $\mathbf{r},\mathbf{s}\in\mathbb{R}^{n}$. Since the accumulation is performed on FP32, the rounding error is ignored here. Therefore, the addition only transfer and accumulate the previous rounding error:
$|H(\mathbf{r}^{T})H(\mathbf{s})-\mathbf{r}^{T}\mathbf{s}|\leq\epsilon_{H}\sum_{i=1}^{n}(|r_{i}|+|s_{i}|).$ (23)
The upper bound of rounding error of the exponent $\beta=\mathbf{u}^{T}\mathbf{v}$ in Section 4 is
$|\hat{\beta}-\beta|\leq S=S_{1}+S_{2}$ (24)
$S_{1}=\epsilon_{H}\sup_{p\in[0,w]\times[0,h]}(1+|p_{x}|+|p_{y}|+|p_{x}^{2}|+|p_{x}p_{y}|+|p_{y}|^{2})$
$S2=\epsilon_{H}\sup_{G^{\prime}\in\mathcal{G}^{\prime}}\sum_{i=0}^{5}|v_{i}(G^{\prime})|,$
where $\mathbf{v}(G^{\prime})$ is the Gaussian vector of projected Gaussian $G^{\prime}$.
Noticing that $S_{2}\geq 0$, therefore we have derived a lower bound of the upper bound $S$:
$S\geq S_{1}\geq\epsilon_{H}\cdot(h^{2}+hw+w^{2}).$ (25)
We can conclude that
$S\geq O(h^{2}+hw+w^{2}).$ (26)
The quadratic term of $h,w$ significantly contributes the rounding error. if taking the screen resolution of 720p, i.e., $1280\times 720$, the upper bound exceeds
$S\geq S_{1}\geq 3\times 10^{3}$ (27)
To compute the relative error, we will find the range of $\beta$, which is not culled:
$\frac{1}{255}\leq\exp(\beta)\leq 1,$ (28)
then $\beta\in[-\ln(255),0]$, then the relative error exceeds:
$\Delta\geq\frac{S}{|-\ln(255)|}\geq\frac{S_{1}}{\ln(255)}>37500\%$ (29)
B.2 Rounding Error on Local Coordinates
If applying local Coordinates, the term $S_{1}$ is
$S_{1}=\epsilon_{H}\sup_{\Delta p\in[-8,8]^{2}}(1+|\Delta p_{x}|+|\Delta p_{y}|+|\Delta p_{x}^{2}|+|\Delta p_{x}\Delta p_{y}|+|\Delta p_{y}|^{2}).$ (30)
Then we can numerically calculate the value of $S_{1}$:
$S_{1}=\epsilon_{H}(1+8+8+8^{2}+8^{2}+8^{2})=0.204=O(1).$ (31)
Then we will analyze the other contributor of the upper bound $S$, $S_{2}$.
At first we find the range of the primitives of projected Gaussians:
Means at local coordinates $\Delta\mu$
: Since the coordinate will be left or upside of the tile center, the coordinate value can be negative. Therefore, $\Delta p\in[-w,w]\times[-h,h]$.
logarithm of opacity $-\log(o)$
: In implementation, the opacity $o<\frac{1}{255}$ will be pruned at the preprocessing stage, therefore $-\log(o)\in[-\ln(255),0]$.
Inverse covariance $\Sigma^{\prime-1}=\begin{bmatrix}a&b\\ b&c\end{bmatrix}$
At first we assume that the major and the minor axis of the ellipse are larger than $0.5$, otherwise it will not be rendered. Then the eigenvalues of the projected covariance $\Sigma^{\prime}$ is greater than $0.5$:
$\lambda_{1},\lambda_{2}\geq 0.5.$ (32)
And the eigenvalues of the inverse covariance $\Sigma^{\prime-1}$ satisfies:
$0<\lambda_{1}^{-1},\lambda_{2}^{-1}\leq 2$ (33)
Consider the characteristic polynomial of $\Sigma^{\prime-1}$:
$f(\lambda)=|\lambda I-\Sigma^{-1}|=\lambda^{2}-(a+c)\lambda+ac-b^{2}.$ (34)
These two eigenvalue are the roots of $f(\lambda)=0$, and:
$\lambda_{1}^{-1}+\lambda_{2}^{-1}\in(0,4],\quad\lambda_{1}^{-1}\lambda_{2}^{-1}\in(0,4].$ (35)
Applying Vieta’s formulas:
$a+c\in(0,4],\quad ac-b^{2}\in(0,4].$ (36)
Since $ac-b^{2}\geq 0$, then $ac\geq b^{2}\geq 0$.Therefore, we can conclude that:
$0<a\leq 4,\quad 0<c\leq 4,$ (37)
and
$b^{2}\leq 4+ac\leq 4+\left(\frac{a+c}{2}\right)^{2}\leq 8,$ (38)
$-2\sqrt{2}\leq b\leq 2\sqrt{2}.$ (39)
To simply the notation, we note $\textbf{v}_{\Delta}=[v_{0},v_{1},v_{2},v_{3},v_{4},v_{5}]^{T}$ Next, we can compute the term $S_{2}$:
$S_{2}$ $=\epsilon_{H}\sup(|v_{0}|+|v_{1}|+|v_{2}|+|v_{3}|+|v_{4}|+|v_{5}|)$ (40)
$=\epsilon_{H}\sup(|v_{0}|)+\epsilon_{H}(2\sqrt{2}+4)(h+w+1)$
$=\epsilon_{H}\sup(|v_{0}|)+O(h+w).$
Therefore, we can conclude that:
$S=S_{1}+S_{2}=\epsilon_{H}\sup(|v_{0}|)+O(h+w).$ (41)
Then we turn to bound the value of $\beta$:
$\beta$ $=v_{0}+\Delta{p_{x}}v_{1}+\Delta p_{y}v_{2}+\Delta p_{x}^{2}v_{3}+\Delta p_{x}\Delta p_{y}v_{4}+\Delta p_{y}^{2}v_{5}$ (42)
$=v_{0}+O(h+w)+O(h+w)+O(1)+O(1)+O(1)$
$=v_{0}+O(h+w)$
Since we only consider the $\beta$ value which contributes the pixel color, $\beta\in[-\ln(255),0]$. Then
$v_{0}=\beta-O(h+w)=O(h+w).$ (43)
Therefore, the upper bound of the of the error of $\beta$ satisfies
$|\hat{\beta}-\beta|\leq S=O(h+w)+\epsilon_{H}O(h+w)=O(h+w).$ (44)
if $\beta\in[-\ln(255),0]$.

### v2

@@ 5.4.2. Training
6. Conclusion
In this paper, we propose TC-GS, a high-performance rendering module for accelerating 3D Gaussian Splatting. TC-GS optimizes the rendering pipeline by decomposing alpha-blending into two components, i.e., Gaussian-related and pixel-related matrices. Through reformulating alpha-blending as matrix multiplications, TC-GS
[…]
Acknowledgements.
This work is supported by the National Key R&\&D Program of China (No. 2022ZD0160201), Shanghai Artificial Intelligence Laboratory, and Shanghai Municipal Science and Technology Major Project.
Appendix A Derivation and Proof of the Upper Bound of Rounding Error on TC-GS
A.1. Floating-Point Numbers
In computing, floating-point numbers arithmetic is arithmetic on subset of real numbers, which can be represented with a finite amount of memory. A floating-point number $r$ is formed by a significand $s$, base $b$ and exponent $e$:
(22) $r=s\cdot b^{e}.$
In this paper, we only consider the binary formats, whose base $b$ is $2$.
In most modern computing systems, a floating-point number is consists of 3 parts:
  • Sign bit: Represents the sign of the number. A sign bit of 0 indicates a positive number, and $1$ indicates a negative number.
  • Exponent: Represents the scale of the number. The exponent is stored with an offset.
  • Significand: Represents the mantissa of the floating-point number. The number of bits of significand is the precision.
For example, if the significand has $m$ bits, the real value of the floating-point number is:
(23) $(-1)^{\text{sign}}(1.b_{m-1}b_{m-2}\dots b_{0})_{2}\times 2^{\text{exponent - offset}},$
where $b_{m-1}b_{m-2}\dots b_{0}$ is the significand bits. Table 7 shows 3 floating-point formats in this paper. FP32 is also known as single-precision floating-point format, and FP16 is called half-precision floating-point format.
Format Sign Exponents Significand Machine Epsilon
FP16 11bit 55bits 1010bits $\varepsilon_{H}=9.77\times 10^{-4}$
TF32 11bit 88bits 1010bits $\varepsilon_{H}=9.77\times 10^{-4}$
FP32 11bit 88bits 2323bits $\varepsilon_{F}=1.19\times 10^{-7}$
Table 7. 3 binary floating-point formats in this paper.
A.2. Machine Epsilon
Rounding is a procedure for representation of a real number $r$ $(r\neq 0)$ into a floating-point format as $R(r)$, where $R$ is the rounding function. The relative error $\varepsilon$ of the rounding procedure is defined as:
(24) $\varepsilon(r)=\dfrac{|R(r)-r|}{|r|}.$
For a fixed floating-point system and a rounding procedure, machine epsilon $\varepsilon_{\text{mach}}$ is an upper bound of relative error if no underflow or overflow occurs:
(25) $|R(r)-r|\leq\varepsilon_{\text{mach}}|r|.$
Proposition A.1.
If $\varepsilon_{\text{mach}}$ is a machine epsilon of a floating-point format, the rounded value of a real number $r$ satisfies:
(26) $R(r)=r+O(\varepsilon_{\text{mach}}).$
Proof.
Since the $\varepsilon_{\text{mach}}$ does not depend on $r$, then Inequality (25) shows $|R(r)-r|=O(\varepsilon_{\text{mach}})$. Therefore, $R(r)-r=O(\varepsilon_{\text{mach}})$.
Proposition A.2.
If a binary floating-point format has $m$ significand bits, $\varepsilon_{\text{mach}}=2^{-m}$ is its machine epsilon.
Proof.
Considering $|r|\in[2^{e},2^{e+1})$ is a real number which does not overflow or underflow in this format, it will lies between two adjacent floating-point numbers, $r_{b}$ and $r_{u}$:
(27) $r_{b}\leq|r|<r_{u}.$
Suppose $x_{b}=(1.b_{m-1}b_{m-2}\dots b_{0})_{2}\times 2^{e}$, then the adjacent $r_{u}$ is:
(28) $r_{u}=(1.b_{m-1}b_{m-2}\dots b_{0})_{2}+(0.00\cdots 1)_{2}\times 2^{e}=r_{b}+2^{e-m}.$
Since $R(r)$ is either $r_{u}$ or $r_{b}$,
(29) $|R(r)-r|\leq|r_{u}-r_{b}|=2^{e-m}=2^{e}\cdot 2^{-m}\leq 2^{-m}|r|.$
Therefore, $2^{-m}$ is a machine epsilon of the floating-point format.
From Proposition A.2, we can conclude that the machine epsilon is only depend on the number of significand bits.
In this paper, we represent the machine epsilon of TF32 and FP16 as $\varepsilon_{H}=9.77\times 10^{-4}$, and the machine epsilon of FP32 as $\varepsilon_{F}=1.19\times 10^{-7}$. Meanwhile, the rounding process of FP16 and TF32 is denoted as $R_{H}$, while the rounding process of FP32 is represented as $R_{F}$.
Furthermore, if the format is fixed, the rounding error is only depend on the absolute value $|r|$. The larger absolute value contribute the larger rounding error. The core idea of G2L in Section 4.3 is reducing the absolute value during the computing process.
A.3. Rounding Error of Dot Product with Tensor Cores
Matrix multiplication can be interpreted as the dot product of vectors at each element, as discussed in Section 4.2. Therefore, to analyze the rounding error in matrix multiplication, it suffices to examine the error in the dot product:
(30) $\beta=\mathbf{u}^{T}\mathbf{v}=\sum_{i=1}^{n}u_{i}v_{i},$
where $n$ is the length of the vectors.
The dot product can be decomposed as a multiplication step and a accumulation step. On Tensor Cores, the multiplication step are performed on half-precision format (FP16 or TF32) and the accumulation step can be executed on FP32 format.
We omit the detailed analysis on terms of small $\varepsilon_{H}^{2}$ and $\varepsilon_{F}$ and represent them as $O(\varepsilon_{H}^{2})$ and $O(\varepsilon_{F})$.
Proposition A.3.
On Tensor Cores, the absolute error of $\beta=\mathbf{u}^{T}\mathbf{v}$ satisfies
(31) $|\hat{\beta}-\beta|\leq 2\varepsilon_{H}\sum_{i=1}^{n}|u_{i}v_{i}|+O(\varepsilon_{H}^{2}+\varepsilon_{F}),$
where $\hat{\beta}$ is the computed value of $\beta$ with Tensor Cores.
Proof.
At first, we consider product of $u_{i}$ and $v_{i}$. Before multiplication, these two real number are stored in half-precision format as $R_{H}(u_{i})$ and $R_{H}(v_{i})$ satisfying
(32) $|R_{H}(u_{i})|\leq|R_{H}(u_{i})-u_{i}|+|u_{i}|\leq|u_{i}|\varepsilon_{H}+|u_{i}|=(1+\varepsilon_{H})|u_{i}|$
and
(33) $|R_{H}(v_{i})|\leq(1+\varepsilon_{H})|v_{i}|.$
Therefore,
(34) $|R_{H}(u_{i})R_{H}(v_{i})|\leq(1+\varepsilon_{H})^{2}(|v_{i}||u_{i}|)=(1+2\varepsilon_{H})|u_{i}v_{i}|+O(\epsilon_{H}^{2}).$
The value product are stored in FP32 format for next accumulation step on Tensor Cores:
(35) $|R_{F}(R_{H}(u_{i})R_{H}(v_{i}))|$ $\leq(1+\varepsilon_{F})|R_{H}(u_{i})R_{H}(v_{i})|$
$\leq(1+\varepsilon_{F})((1+2\varepsilon_{H})|u_{i}v_{i}|+O(\varepsilon_{H}^{2}))$
$\leq(1+2\varepsilon_{H})|u_{i}v_{i}|+O(\varepsilon_{H}^{2})+O(\varepsilon_{F}).$
Notice that the rounding function does not change the sign of the number. Therefore, relative error of the product is:
(36) $\varepsilon(u_{i}v_{i})$ $=\dfrac{|u_{i}v_{i}-R_{F}(R_{H}(u_{i})R_{H}(v_{i}))|}{|u_{i}v_{i}|}$
$=\dfrac{||u_{i}v_{i}|-|R_{F}(R_{H}(u_{i})R_{H}(v_{i}))||}{|u_{i}v_{i}|}$
$\leq 2\varepsilon_{H}+O(\varepsilon_{H}^{2})+O(\varepsilon_{F}).$
To simply the proof, we represent $a_{i}$ as the accumulator on Tensor Cores:
(37) $a_{i}=R_{F}(R_{H}(u_{i})R_{H}(v_{i}))=u_{i}v_{i}(1\pm\varepsilon(u_{i}v_{i}))$
We assume that the addition is performed from left to right. Since the sum of each two values are stored in FP32 as a new accumulator, the computed cumulative sum $s_{i}$ of $\{a_{i}\}_{i=1}^{n}$ is:
(38) $s_{1}=R_{F}(a_{1})=a_{1},\quad s_{i}=R_{F}(s_{i-1}+a_{i}).$
Therefore, the computed value of $\beta$ on Tensor Cores is:
(39) $\hat{\beta}=R_{F}(R_{F}(\cdots R_{F}(R_{F}(a_{1}+a_{2})+a_{3})\cdots)+a_{n})=s_{n}.$
Next, we will prove that $s_{i}=\sum_{j=1}^{i}a_{j}+O(\varepsilon_{F})$ inductively:
  1. (1)
    It is obvious that $s_{1}=a_{1}=a_{1}+O(\varepsilon_{F})$;
  2. (2)
    For a positive integer $i<n$, we assume that $s_{i}=\sum_{j=1}^{i}a_{j}+O(\varepsilon_{F})$. According to Proposition 26,
    (40) $s_{i+1}=R_{F}(s_{i}+a_{i})=s_{i}+a_{i}+O(\varepsilon_{F})=\sum_{j=1}^{i+1}a_{j}+O(\varepsilon_{F}).$
According to Equation (37),
(41) $\hat{\beta}=s_{n}=\sum_{i=1}^{n}a_{i}+O(\varepsilon_{H})=\sum_{i=1}^{n}(u_{i}v_{i}\pm\varepsilon(u_{i}v_{i}))+O(\varepsilon_{F}).$
To sum up, the absolute error of $\beta$ satisfies
(42) $|\hat{\beta}-\beta|$ $=\left|\sum_{i=1}^{n}u_{i}v_{i}(1\pm\varepsilon(u_{i}v_{i}))+O(\varepsilon_{F})-\sum_{i=1}^{n}u_{i}v_{i}\right|$
$=\left|\sum_{i=1}^{n}\pm u_{i}v_{i}\varepsilon(u_{i}v_{i})+O(\varepsilon_{F})\right|$
$=\left|\sum_{i=1}^{n}\pm u_{i}v_{i}(2\varepsilon_{H}+O(\varepsilon_{H}^{2})+O(\varepsilon_{F}))+O(\varepsilon_{F})\right|$
$=\left|\sum_{i=1}^{n}\pm 2u_{i}v_{i}\varepsilon_{H}+O(\varepsilon_{H}^{2})+O(\varepsilon_{F})\right|$
$\leq 2\varepsilon_{H}\sum_{i=1}^{n}|u_{i}v_{i}|+O(\varepsilon_{H}^{2}+\varepsilon_{F}).$
The proof of Proposition A.3 is completed here.
A.4. Absolute Error of $\beta$ on Global Coordinates
This section discusses the absolute of exponent $\beta$ on global coordinates.
The absolute error of $\beta$ is dependent on $\mathbf{v}$, and $\mathbf{v}$ is depend on the parameters of projected Gaussian $G^{\prime}$. Therefore, we discusses the range of Gaussians’ parameters at first.
We represent the projected Gaussian as $G^{\prime}=(\mu^{\prime},\Sigma^{\prime},o,c)$ with $\mu^{\prime}=(\mu^{\prime}_{x},\mu^{\prime}_{y})$ and symmetric $\Sigma^{\prime-1}=\begin{bmatrix}\sigma_{11}&\sigma_{12}\\ \sigma_{21}&\sigma_{22}\end{bmatrix}$ satisfying $\sigma_{12}=\sigma_{21}$. The alpha value at pixel $p=(p_{x},p_{y})$ is $\alpha=\mathrm{e}^{\beta}=\mathrm{e^{\mathbf{u^{T}}\mathbf{v}}}$ as Section 4.2 discussed.
The width and height of rendered image are represented as $w$ and $h$ respectively.
Means $\mu^{\prime}$:
To simplify the analysis, we only consider Gaussians whose center within the screen space. Therefore, $\mu^{\prime}\in[0,w]\times[0,h]$.
Opacity $o$:
The fragment with alpha value below $\frac{1}{255}$ will be culled. Since $\alpha\leq o$ since $\Sigma^{\prime-1}$ is semi-definite, $o\in[\frac{1}{255},1]$.
Inverse Covariance $\Sigma^{\prime-1}$:
The covariance $\Sigma^{\prime}$ identifies the shape of the ellipse on the screen space. If $\lambda_{1}$ and $\lambda_{2}$ are the two eigenvalues of $\Sigma^{\prime}$, then length of two semi-axes of the ellipse are $\lambda_{1}$ and $\lambda_{2}$.
We assume the length of semi-axis is larger than 0.5, otherwise the Gaussian will not be rendered:
(43) $\lambda_{1},\lambda_{2}\geq 0.5.$
Then the eigenvalues of the inverse covariance $\Sigma^{\prime-1}$ satisfies:
(44) $0<\lambda_{1}^{-1}+\lambda_{2}^{-1}\leq 4,\quad 0<\lambda_{1}^{-1}\lambda_{2}^{-1}\leq 4.$
Consider the characteristic polynomial of $\Sigma^{\prime-1}$:
(45) $f(\lambda^{-1})=|\lambda^{-1}I-\Sigma^{-1}|=\lambda^{-2}-(\sigma_{11}+\sigma_{22})\lambda^{-1}+\sigma_{11}\sigma_{22}-\sigma_{12}^{2}.$
These two eigenvalues satisfy $f(\lambda^{-1})=0$. Applying the Vieta’s Formulas,
(46) $0<\sigma_{11}+\sigma_{22}\leq 4,\quad 0<\sigma_{11}\sigma_{22}-\sigma_{12}^{2}\leq 4.$
Then,
(47) $\sigma_{12}^{2}<\sigma_{11}\sigma_{22}\leq\left(\frac{\sigma_{11}+\sigma_{22}}{2}\right)^{2}\leq 4.$
Therefore,
(48) $0<\sigma_{11}<4,\quad-2\leq\sigma_{12}\leq 2,\quad 0<\sigma_{22}<4.$
According to Proposition A.3, the absolute error of $\beta$ on Frag2Mat is:
(49) $|\hat{\beta}-\beta|\leq 2\varepsilon_{H}\sum_{i=0}^{5}|u_{i}v_{i}|+O(\varepsilon_{H}^{2}+\varepsilon_{F}).$
Applying the global coordinate, the coordinate of pixel satisfies $p=(p_{x},p_{y})\in[0,w]\times[0,h]$.
Then,
(50) $|u_{0}v_{0}|=\left|\ln(o)-\frac{1}{2}\sigma_{11}\mu_{x}^{2}-\sigma_{12}\mu_{x}\mu_{y}-\sigma_{22}\mu_{y}^{2}\right|=O(w^{2}+wh+h^{2}),$
(51) $|u_{1}v_{1}|=|p_{x}(\sigma_{11}\mu_{x}+\sigma_{12}\mu_{y})|=O(w^{2}+wh),$
(52) $|u_{2}v_{2}|=|p_{y}(\sigma_{12}\mu_{x}+\sigma_{22}\mu_{y})|=O(wh+h^{2}),$
(53) $|u_{3}v_{3}|=\left|p_{x}^{2}(-\frac{1}{2}\sigma_{11})\right|=O(w^{2}),$
(54) $|u_{4}v_{4}|=|p_{x}p_{y}(-\sigma_{12})|=O(wh),$
(55) $|u_{5}v_{5}|=\left|p_{y}^{2}(-\frac{1}{2}\sigma_{22})\right|=O(h^{2}).$
Therefore, the absolute error of $\beta$ satisfies quadratic
(56) $|\hat{\beta}-\beta|=O(w^{2}+wh+h^{2})\varepsilon_{H}+O(\varepsilon_{H}^{2}+\varepsilon_{F}).$
A.5. Absolute Error of $\beta$ on Local Coordinates
This section discusses the absolute of exponent $\beta$ on local coordinates.
At first, we discuss the range of Gaussians’ parameters as Section A.4. In this section, $p(\mathcal{T})$ is the center of tile $\mathcal{T}$, and the size of each tile is $16\times 16$.
Means $\Delta\mu^{\prime}$:
Since $\mu^{\prime},p(\mathcal{T})\in[0,w]\times[0,h]$, $\Delta\mu^{\prime}=\mu^{\prime}-p(\mathcal{T})\in[-w,w]\times[-h,h]$.
Opacity $o$:
The opacity is the same as Gaussians on global coordinates. $o\in[\frac{1}{255},1]$.
Inverse Covariance $\Sigma^{\prime-1}$:
. The covariance is the same as Gaussians on global coordinates: $0<\sigma_{11},\sigma_{12}<4,-2\leq\sigma_{12}\leq 2$.
All pixels on local coordinates is within a tile, then $\Delta p=(\Delta p_{x},\Delta p_{y})\in[-8,8]^{2}$.
Therefore,
(57) $|u_{1}v_{1}|=|\Delta p_{x}(\sigma_{11}\Delta\mu_{x}+\sigma_{12}\Delta\mu_{y})|=O(w+h),$
(58) $|u_{2}v_{2}|=|\Delta p_{y}(\sigma_{12}\Delta\mu_{x}+\sigma_{22}\Delta\mu_{y})|=O(w+h),$
(59) $|u_{3}v_{3}|=\left|\Delta p_{x}^{2}(-\frac{1}{2}\sigma_{11})\right|=O(1),$
(60) $|u_{4}v_{4}|=|\Delta p_{x}\Delta p_{y}(-\sigma_{12})|=O(1),$
(61) $|u_{5}v_{5}|=\left|\Delta p_{y}^{2}(-\frac{1}{2}\sigma_{22})\right|=O(1).$
According to Proposition A.3,
(62) $|\hat{\beta}-\beta|\leq(2|u_{0}v_{0}|+O(w+h))\varepsilon_{H}+O(\varepsilon_{H}^{2}+\varepsilon_{F}).$
We only care the precision of fragments which are not culled, then $\ln(255)\leq\beta<0$. Since $\beta=\sum_{i=0}^{5}u_{i}v_{i}$, then $u_{0}v_{0}=\beta-\sum_{i=1}^{5}u_{i}v_{i}$. Therefore,
(63) $|u_{0}v_{0}|\leq|\beta|+\sum_{i=1}^{5}|u_{i}v_{i}|=O(w+h).$
In conclusion, the absolute error of $\beta$ satisfies
(64) $|\hat{\beta}-\beta|\leq O(w+h)\varepsilon_{H}+O(\varepsilon_{H}^{2}+\varepsilon_{F}).$
Appendix B Notations
@@
$c$ Color of 3D Gaussian $W$ Matrix projecting 3D position into 2D plane
$J$ Jacobian of projection $\mu^{\prime}=(\mu^{\prime}_{x},\mu^{\prime}_{y})$ Mean of projected Gaussian
$\Sigma^{\prime}=\begin{bmatrix}\sigma_{11}&\sigma_{12}\\ \sigma_{21}&\sigma_{22}\end{bmatrix}^{-1}$ Covariance of projected Gaussian $G^{\prime}$ Projected Gaussian
$\mathcal{G}^{\prime}$ Set of projected Gaussians $p=(p_{x},p_{y})$ Pixel
@@
$C$ Rendered pixel color $\alpha$ Alpha value of fragments.
@@
NN Number of splats. $\mathcal{E}$ Complexity of the pipeline in 3DGS
@@
$\Delta p=(\Delta p_{x},\Delta p_{y}).$ Pixel on local coordinates $\Delta\mu^{\prime}=(\Delta\mu_{x}^{\prime},\Delta\mu_{y}^{\prime})$ Projected mean on local coordinates
$\varepsilon_{H}$ Machine epsilon of FP16 or TF32 $\varepsilon_{F}$ Machine epsilon of FP32
$R_{H}$ Rounding function of FP16 or TF32 $R_{F}$ Rounding function of FP32
$\hat{\beta}$ Computed value of $\beta$ with Tensor Cores $\lambda_{1},\lambda_{2}$ Two eigenvalues of $\Sigma^{\prime}$
Table 8. Notations in this paper.